OpenStudy (anonymous):

Graph the function. Identify the axis of symmetry and the vertex f(x)=x^2-4x+9

OpenStudy (anonymous):

y=x^2-4x+9 y=(x^2-4x+4)+5 y=(x-2)^2+5 then vertex is (2,5) and axis of symmetry is x=2

OpenStudy (anonymous):

thanks and thank you for explaining how to get the answer.

OpenStudy (anonymous):

a simple way for finding x of vertex is it to use this formula x=-b/2a x=4/2=2 then put 2 in the function to find the y of vertex f(2)=4-8+9=5 so vertex is (2,5) and axis of symmetry is x=-b/2a